A ray of light is incident on a plane surface of refractive index√3 at a certain angle it is found that the reflected and refracted rays are perpendicular to each other then the angle of incidence is ?
1
Expert's answer
2020-05-25T10:55:22-0400

Let the angle of incidence, angle of reflection, and angle of refraction be i , r and r′, respectively.

Now, as per the equation

"(90^\u2218\u2212r)+(90^\u2218\u2212r^\u2032)=90^\u2218"

"r^\u2032=(90^\u2218\u2212i)"  (because i=r, in case of reflection according to Snell's law, 1 sin i =μsinr′)

"or\ \sin i= μ \sin r^′"

"\sin i =μ\sin(90∘−i)⇒\tan i=μ"

"i=\tan^{−1}μ=tan^{−1}√3=60^∘"